# Emerging Activity Temporal Hypergraph (EATH), a model for generating realistic time-varying hypergraphs

### v1

@@ II.1 Definitions and notations
We consider a time-varying hypergraph $\mathscr{H}$ over the time interval $(0,T]$ and we represent it as a sequence of $T/\delta t$ static hypergraphs =$\mathscr{H}=\{\mathcal{H}_{t}\}_{t=1}^{T/\delta t}$, with time-resolution $\delta t$. At each time $t$ the system is represented by the unweighted instantaneous hypergraph $\mathcal{H}_{t}=(\mathcal{V}_{t},\mathcal{E}_{t})$, with $N_{t}=|\mathcal{V}_{t}|$ active nodes participating in $E_{t}=|\mathcal{E}_{t}|$ hyperedges of arbitrary size $m\in[2,M]$ (we define a node as “active” if it is involved in at least one hyperedge of size $\geq 2$, i.e., when it is not isolated). The nodes and hyperedges within $\mathcal{H}_{t}$ are the ones that have been active at least once in the time interval $((t-1)\delta t,t\delta t]$ [12, 10].
@@
To each node $i$ we can finally assign its aggregated properties in the aggregated hypergraph and in its pairwise projection: its hyperdegree $D(i)$, its hyper-strength $S(i)=\sum_{e\in\mathcal{E}|i\in e}w_{e}$ in $\mathcal{H}$, and its degree $D_{proj}(i)$ and its strength $S_{proj}(i)=\sum_{l\in\mathcal{L}|i\in l}w_{l}$ in $\mathcal{G}$.
@@ II.2 Empirical temporal hypergraphs
Real time-varying networks and hypergraphs are characterized by non-trivial temporal and higher-order properties [21, 9, 11]. Here, we consider eight different temporal hypergraphs of face-to-face interactions between individuals, collected in various contexts (see Methods for their detailed description): schools in France and in the USA (“LyonSchool”, “Thiers13”, “Utah_elem”, “Utah_mid”), hospital (“LH10”), and conferences (“SFHH”, “ICCSS17”, “WS16”). For all these datasets the entire system’s activity undergoes important variations in time, as shown in Fig. 1a, where we report an example of the temporal evolution of the total number of interactions $E_{t}$ (see the Supplementary Material, SM, for the other datasets). Periods of high activity are followed by intervals of inactivity, and different orders of interactions can be more or less active at different times (see SM). Different orders are also characterized by different frequencies of appearance: small interaction sizes are encountered more often, and large hyperedges are much less frequent, as shown by Fig. 1b where the hyperedges size distributions $\Psi(m)$ are reported for all the datasets. These observations provide a first insight of the complexity of empirical hypergraphs and will constitute the basis on which we will start to build our generation model.
Figure 1c-h displays the distribution of six quantities characterizing the temporal activation of nodes and of hyperedges. Starting with the point of view of nodes, Fig. 1c reports the distribution of node event durations $P(T)$ for all the nodes of the hypergraph. We recall that an event for node $i$ starts at time $t$ if $i$ goes from being isolated at $t-1$ to being involved in at least one hyperedge at $t$, and the event ends when $i$ becomes again isolated (when it is no longer involved in any hyperedge). Note that an event for node $i$ is considered to go on even if the hyperedges in which it is involved during the event change: the event describes a temporal interval between two times in which $i$ is isolated. The duration of a node event hence corresponds to the number of consecutive time steps in which the node is active. Figure 1c shows that all the considered hypergraphs are characterized by heavy-tailed distributions of the node event durations.
@@
Figure 2: Nodes and hyperedges temporal heterogeneity. For each node and hyperedge we compute the burstiness values $\Delta B_{\tau}$ and $\Delta B_{T}$ of their event and inter-event duration distributions. In each panel, each point corresponds to a node (a,c,e,g,i,k,m,o) or a hyperedge (b,d,f,h,j,l,n,p) and we show the scatterplot between $\Delta B_{\tau}$ and $\Delta B_{T}$. The gray dashed lines correspond to the reference $\Delta B=0$. Panels q, r show the distributions $P(\Delta B_{\tau})$ and $P(\Delta B_{T})$, respectively for nodes and hyperedges, for the SFHH dataset. We consider only the nodes and hyperedges with at least $10$ activation events.
@@
The distributions shown in Fig. 1 are built considering all nodes or all hyperedges. These broad distributions could however result from a superposition of either narrow or broad distributions of durations at the single node or hyperedge level, and the dynamical properties of different nodes and hyperedges could potentially be either homogeneous or heterogeneous. We thus investigate the dynamical properties of individual nodes and hyperedges by computing the burstiness of the distribution of event and inter-event durations of each node and hyperedge: Fig. 2 reports for each dataset a scatterplot of the resulting burstiness values ($\Delta B_{T}$ vs. $\Delta B_{\tau}$). Figure 2q, r report moreover the distributions of the node and hyperedge burstiness values for the SFHH conference dataset [49, 50, 51, 50] (see the SM for other datasets). Interestingly, these distributions span a broad interval of values (almost all positive): this reveals a clear heterogeneity of nodes and hyperedges temporal behaviors, some exhibiting a very bursty activity while other have statistics of event or inter-event durations closer to a Poissonian dynamics. Moreover, the scatterplots do not exhibit any correlation between the burstiness value for the event durations and the one for the inter-event durations.
@@
While the statistics measured above do not distinguish between hyperedges of different sizes, Figure 1b shows that not all sizes are equally probable in empirical hypergraphs. As Fig. 1b displays a global distribution, we also investigate whether different nodes display different behaviors in terms of the sizes of the hyperedges in which they participate. To this aim, we consider the total time spent by each node $i$ in hyperedges of order $m$, ${T}_{i}^{m}=\sum_{e\in\mathcal{F}(i)||e|=m}w_{e}$, with $\mathcal{F}(i)$ the set of hyperedges in which $i$ is involved in $\mathcal{H}$. We then rank the nodes according to this measure for each order $m$ and we compare the rankings at different orders: Fig.  3 gives the Pearson’s correlation coefficient $\rho(m,m^{\prime})$ between the rankings at orders $m$ and $m^{\prime}$. In general, adjacent orders tend to have similar node rankings, but in some cases the matrices are organized in blocks, with very low correlation values between orders belonging to different blocks. For instance, in the SFHH dataset the node rankings at orders between 2 and 4 are very different from those at orders between 6 and 9. Thus, the nodes that are the most important for large interactions are not the most present in small interactions, and vice-versa, revealing a further heterogeneity among nodes, which can be divided into categories depending on how they allocate their social activity in groups of different sizes. Note that these differences in behaviour emerge from comparing the rankings at different group sizes, since the nodes allocate differently their activity at the various orders; however, for all nodes the largest $T_{i}^{m}$ are obtained for small sizes $m$ (see SM).
@@ III Emerging Activity Temporal Hypergraph (EATH) model
We propose a model to generate a synthetic time-varying hypergraph =$\mathscr{H}=\{\mathcal{H}_{t}\}_{t=1}^{T/\delta t}$ over the time interval $(0,T]$, with time-resolution $\delta t$, with hyperedges of size $m\in[2,M]$ and with a population of $N$ nodes.
@@ III.1 Nodes and system activity dynamics
where $\gamma\in[0,1]$, $h(t)$ and $l(t)$ are respectively the set of nodes in the high- and low-activity phase at time $t$. Hereafter we will fix $\gamma=10^{-3}$ (see SM for other values).
@@
$r_{l\to h}(i,t)=\Lambda_{t}\varrho_{l}\frac{a_{T}(i)}{\langle a_{T}(i)\rangle},$ (2)
@@
where $\Lambda_{t}\in[0,1]$ is the external modulation, which impacts all nodes transition probabilities in the same way; the parameters $\varrho_{l}$ and $\varrho_{h}$ set the time-scales of the evolution, i.e. the average time nodes spend in the low-activity state $\Delta_{l}=\langle\frac{1}{r_{l\to h}(i,t)}\rangle$, and in the high-activity state $\Delta_{h}=\langle\frac{1}{r_{h\to l}(i,t)}\rangle$. Note that nodes with higher persistence activity $a_{T}(i)$ spend more time in the high-activity state as they have a larger ratio $r_{l\to h}(i,t)/r_{h\to l}(i,t)$.
@@
$E_{t}=\lambda\frac{A_{t}}{\langle A_{t}\rangle},$ (4)
where $A_{t}$ is the cumulative contribution of all nodes’ activities, $A_{t}=\sum_{i\in\mathcal{V}}a_{t}(i)$, and the parameter $\lambda$ determines the average number of hyperedges generated (as $\langle E_{t}\rangle=\lambda$). The sizes of the $E_{t}$ active hyperedges at time $t$ are distributed according to $\Psi(m)$ as described below.
@@ III.2 Creation of the active hyperedges at each snapshot
  • with probability $p$, $e$ is generated from scratch randomly: this models the generation of new groups not related to interactions active at the previous time $t-1$. The nodes forming $e$ are selected with probability proportional to their activity $a_{t}(i)$, to their order propensity $\varphi_{i}(m)$ and to the memory $\tilde{\omega}_{0}$ within the group (see Methods for details).
  • with probability $(1-p)$, a short-term memory mechanism is instead used, namely, the hyperedge is generated as the continuation of a hyperedge $e^{\prime}$ that was active at $t-1$, and had size $m^{\prime}$ equal to $m-1$, $m$ or $m+1$. $e$ will then be either equal to $e^{\prime}$ (if $m^{\prime}=m$), or consist in a group $e^{\prime}$ gaining (if $m^{\prime}=m-1$) or losing (if $m^{\prime}=m+1$) a member. This allows to model aggregation and disaggregation mechanisms. Among all the possible processes $e^{\prime}\to e$ (corresponding each to a hyperedge of $\mathcal{E}_{t-1}$ with size $\in\{m-1,m,m+1\}$), the actual one is chosen with probability proportional to the activities $a_{t}(i)$ of $e$’s nodes, to their order propensities $\varphi_{i}(m)$, to the long-term memory $\tilde{\omega}_{0}$ within $e$, and to the duration of the hyperedge $e^{\prime}$. See the Methods for more details. Note that here we consider that only one node can leave/join an existing group at each time step, however more complex splitting and merging mechanisms could potentially be considered [11].
@@ III.3 EATH model parameters
The system parameters include: the average number of interactions $\langle E_{t}\rangle$ and their size distribution $\Psi(m)$, defining the connectivity and the type of group activity; the modulation $\Lambda_{t}$ and the time scales $T,\Delta_{h},\Delta_{l}$, which describe temporal patterns due to specific schedules; the memory matrix heterogeneity $\langle\omega_{0}^{2}\rangle/\langle\omega_{0}\rangle^{2}$ and the probability $p$ that determine the balance between the generation of new hyperedges and the reinforcement of already activated groups. The population properties are: the number of nodes $N$, which defines the size of the population; the distribution of activities $a_{h}(i)$, $a_{T}(i)$, and of the order propensity $\varphi_{i}(m)$, which describe the agents’ heterogeneity. In the next Section we describe how to extract these parameters from a dataset in order to generate a surrogate temporal hypergraph of that dataset.
@@ IV.1 From data to model parameters
First, several parameters are simple quantities that can be straightforwardly extracted from the datasets, such as the population size $N$, the duration $T$ and the average number of interactions $\langle E_{t}\rangle$.
@@
  • the probability of generating a new hyperedge $p$ is determined by counting at each time-step $t$ the number of hyperedges $O_{t}$ that already existed at time $t-1$ or that have lost/gained one node with respect to time $t-1$:
    $p=\sum_{t=1}^{T/\delta t}(E_{t}-O_{t})/\sum_{t=1}^{T/\delta t}E_{t}.$ (8)
@@
In the following subsections, we show the results of the generation of surrogate temporal hypergraphs when using the SFHH dataset as original data [49, 50, 51, 50]: this dataset describes face-to-face interactions occurring during a conference with 403 participants, collected using wearable proximity sensors by the SocioPatterns collaboration over a time span of two days, with a temporal resolution of 20 seconds (see Methods for details). In the SM we show analogous results for all the other datasets. Moreover, in addition to hypergraphs generated with the EATH model described above, we also consider a baseline version in which no memory effect is taken into account in the generation process (EATHw version): this will allow us to test the importance of the memory mechanisms in reproducing empirical properties.

### v2

# Emerging Activity Temporal Hypergraph: a model for generating realistic time-varying hypergraphs

@@ II.1 Definitions and notations
We consider a time-varying hypergraph $\mathscr{H}$ over the time interval $(0,\mathcal{T}]$ and we represent it as a sequence of $\mathcal{T}/\delta t$ static hypergraphs $\mathscr{H}=\{\mathcal{H}_{t}\}_{t=1}^{\mathcal{T}/\delta t}$, with time-resolution $\delta t$ (the index $t$ goes from 1 to $\mathcal{T}/\delta t$, indicating each static snapshot starting from the first time-window at $t=1$). At each time $t$ the system is represented by the unweighted instantaneous hypergraph $\mathcal{H}_{t}=(\mathcal{V}_{t},\mathcal{E}_{t})$, with $N_{t}=|\mathcal{V}_{t}|$ active nodes participating in $E_{t}=|\mathcal{E}_{t}|$ hyperedges of arbitrary size $m\in[2,M]$ (we define a node as “active” if it is involved in at least one hyperedge of size $\geq 2$, i.e., when it is not isolated). The nodes and hyperedges within $\mathcal{H}_{t}$ are the ones that have been active at least once in the time interval $((t-1)\delta t,t\delta t]$ [12, 10].
@@
To each node $i$ we can finally assign its aggregated properties in the aggregated hypergraph and in its pairwise projection: namely, its hyperdegree $D(i)$, i.e., the number of distinct hyperedges in which $i$ is involved in $\mathcal{H}$, its hyper-strength $S(i)=\sum_{e\in\mathcal{E}|i\in e}w_{e}$, its degree $D_{proj}(i)$, i.e., the number of distinct links in which $i$ is involved in $\mathcal{G}$, and its strength $S_{proj}(i)=\sum_{l\in\mathcal{L}|i\in l}w_{l}$.
@@ II.2 Empirical temporal hypergraphs
Real time-varying networks and hypergraphs are characterized by non-trivial temporal and higher-order properties [21, 9, 11]. Here, we consider eight different temporal hypergraphs of face-to-face interactions between individuals, collected in various contexts (see Methods for their detailed description): schools in France and in the USA (“LyonSchool”, “Thiers13”, “Utah_elem”, “Utah_mid”), hospital (“LH10”), and conferences (“SFHH”, “ICCSS17”, “WS16”). For all these datasets the entire system’s activity undergoes important variations in time, as shown in Fig. 1a, where we report an example of the temporal evolution of the total number of interactions $E_{t}$ (see the Sec. IA of the Supplementary Material, SM, for the other datasets). Periods of high activity are followed by intervals of inactivity, and different orders of interactions can be more or less active at different times (see SM, Sec. IA). Different orders are also characterized by different frequencies of appearance: small interaction sizes are encountered more often, and large hyperedges are much less frequent, as shown by Fig. 1b where the hyperedges size distributions $\Psi(m)$ are reported for all the datasets. These observations provide a first insight of the complexity of empirical hypergraphs and will constitute the basis on which we will start to build our generation model.
Figure 1c-h displays the distribution of six quantities characterizing the temporal activation of nodes and of hyperedges. Starting with the point of view of nodes, Fig. 1c reports the distribution of node event durations $P(T)$ for all the nodes of the hypergraph. We recall that an event for node $i$ starts at time $t$ if $i$ goes from being isolated at t1t-1 to being involved in at least one hyperedge at $t$, and the event ends when $i$ becomes again isolated (when it is no longer involved in any hyperedge). Note that an event for node $i$ is considered to go on even if the hyperedges in which it is involved during the event change: the event describes a temporal interval between two times in which $i$ is isolated. The duration of a node event hence corresponds to the number of consecutive time steps in which the node is active. Figure 1c shows that all the considered hypergraphs are characterized by heavy-tailed distributions of the node event durations.
@@
Figure 2: Nodes and hyperedges temporal heterogeneity. For each node and hyperedge we compute the burstiness values $\Delta B_{\tau}$ and $\Delta B_{T}$ of their event and inter-event duration distributions. In each panel, each point corresponds to a node (a,c,e,g,i,k,m,o) or a hyperedge (b,d,f,h,j,l,n,p) and we show the scatterplot between $\Delta B_{\tau}$ and $\Delta B_{T}$. The gray dashed lines correspond to the reference $\Delta B=0$. Panels q, r show the distributions $P(\Delta B_{\tau})$ and $P(\Delta B_{T})$, respectively for nodes and hyperedges, for the SFHH dataset. We consider only the nodes and hyperedges with at least 1010 activation events.
@@
The distributions shown in Fig. 1 are built considering all nodes or all hyperedges. These broad distributions could however result from a superposition of either narrow or broad distributions of durations at the single node or hyperedge level, and the dynamical properties of different nodes and hyperedges could potentially be either homogeneous or heterogeneous. We thus investigate the dynamical properties of individual nodes and hyperedges by computing the burstiness of the distribution of event and inter-event durations of each node and hyperedge. Fig. 2 reports for each dataset a scatterplot of the resulting burstiness values ($\Delta B_{T}$ vs. $\Delta B_{\tau}$): these plots highlight the complex temporal behavior of nodes (resp. groups) by illustrating simultaneously: (i) the heterogeneity in the nodes (resp. hyperedges) behaviour; (ii) the correlation (or absence thereof) between the two burstiness measures. Figure 2q, r report moreover the distributions of the node and hyperedge burstiness values for the SFHH conference dataset [49, 50, 51, 50] (see the SM, Sec. IC, for other datasets). Interestingly, these distributions span a broad interval of values (almost all positive): this reveals a clear heterogeneity of nodes and hyperedges temporal behaviors, some exhibiting a very bursty activity while other have statistics of event or inter-event durations closer to a Poissonian dynamics. Moreover, the scatterplots do not exhibit any correlation between the burstiness value for the event durations and the one for the inter-event durations: this suggests that the systems are characterized by the presence of nodes and hyperedges with very different temporal behaviour, ranging from very regular statistics in the durations and inter-event times (with $\Delta B_{T}\lesssim 0$ and $\Delta B_{\tau}\lesssim 0$) to very irregular dynamics (with both $\Delta B_{T}$ and $\Delta B_{\tau}$ taking large values), passing through all intermediate combinations of behaviors.
@@
While the statistics measured above do not distinguish between hyperedges of different sizes, Figure 1b shows that not all sizes are equally probable in empirical hypergraphs. As Fig. 1b displays a global distribution, we also investigate whether different nodes display different behaviors in terms of the sizes of the hyperedges in which they participate. To this aim, we consider the total time spent by each node $i$ in hyperedges of order $m$, ${T}_{i}^{m}=\sum_{e\in\mathcal{F}(i)||e|=m}w_{e}$, with $\mathcal{F}(i)$ the set of hyperedges in which $i$ is involved in $\mathcal{H}$. We then rank the nodes according to this measure for each order $m$ and we compare the rankings at different orders: Fig. 3 gives the Pearson’s correlation coefficient $\rho(m,m^{\prime})$ between the rankings at orders $m$ and $m^{\prime}$. In general, adjacent orders tend to have similar node rankings, but in some cases the matrices are organized in blocks, with very low correlation values between orders belonging to different blocks. For instance, in the SFHH dataset the node rankings at orders between 2 and 4 are very different from those at orders between 6 and 9. Thus, the nodes that are the most important for large interactions are not the most present in small interactions, and vice-versa, revealing a further heterogeneity among nodes, which can be divided into categories depending on how they allocate their social activity in groups of different sizes. Note that these differences in behaviour emerge from comparing the rankings at different group sizes, since the nodes allocate differently their activity at the various orders; however, for all nodes the largest $T_{i}^{m}$ are obtained for small sizes $m$ (see SM, Sec. ID).
@@ III Emerging Activity Temporal Hypergraph (EATH) model
We propose a model to generate a synthetic time-varying hypergraph $\mathscr{H}=\{\mathcal{H}_{t}\}_{t=1}^{\mathcal{T}/\delta t}$ over the time interval $(0,\mathcal{T}]$, with time-resolution $\delta t$, with hyperedges of size $m\in[2,M]$ and with a population of $N$ nodes.
@@ III.1 Nodes and system activity dynamics
where $\gamma\in[0,1]$, $h(t)$ and $l(t)$ are respectively the set of nodes in the high- and low-activity phase at time $t$. Hereafter we will fix $\gamma=10^{-3}$ (see SM, Sec. III, for other values).
@@
$r_{l\to h}(i,t)=\Lambda_{t}\varrho_{l}\frac{a_{T}(i)}{\langle a_{T}\rangle},$ (2)
where the average $\langle a_{T}\rangle$ refers to the average of $a_{T}(j)$ over all nodes $j$, and
@@
where $\Lambda_{t}\in[0,1]$ is the external modulation, which impacts all nodes transition probabilities in the same way; the parameters $\varrho_{l}$ and $\varrho_{h}$ set the time-scales of the evolution. Indeed, the average time nodes spend in the low-activity state is $\Delta_{l}=\left\langle\frac{1}{r_{l\to h}(i,t)}\right\rangle$, and the average time they spend in the high-activity state is $\Delta_{h}=\left\langle\frac{1}{r_{h\to l}(i,t)}\right\rangle$ (averages over both nodes and time). Once $\Delta_{l}$ and $\Delta_{h}$ are extracted from the empirical data (or tuned synthetically to desired values), the equations (2) and (3) allow us to obtain the values of the coefficients $\varrho_{l}$ and $\varrho_{h}$ needed to reproduce the desired time-scales (see Methods for more details). Note that nodes with higher persistence activity $a_{T}(i)$ spend more time in the high-activity state as they have a larger ratio $r_{l\to h}(i,t)/r_{h\to l}(i,t)$.
@@
$E_{t}={\xi}\frac{A_{t}}{\langle A_{t}\rangle},$ (4)
where $A_{t}$ is the cumulative contribution of all nodes’ activities, $A_{t}=\sum_{i\in\mathcal{V}}a_{t}(i)$, and the parameter $\xi$} determines the average number of hyperedges generated (as $\langle E_{t}\rangle={\xi}$). The sizes of the $E_{t}$ active hyperedges at time $t$ are distributed according to $\Psi(m)$ as described below.
@@ III.2 Creation of the active hyperedges at each snapshot
  • with probability $p$, $e$ is generated from scratch randomly: this models the generation of new groups not related to interactions active at the previous time t1t-1. The nodes forming $e$ are selected with probability proportional to their activity $a_{t}(i)$, to their order propensity $\varphi_{i}(m)$ and to the memory $\tilde{\omega}_{0}$ within the group (see Methods for details).
  • with probability (1p)(1-p), a short-term memory mechanism is instead used, namely, the hyperedge is generated as the continuation of a hyperedge $e^{\prime}$ that was active at t1t-1, and had size $m^{\prime}$ equal to m1m-1, $m$ or $m+1$. $e$ will then be either equal to $e^{\prime}$ (if $m^{\prime}=m$), or consist in a group $e^{\prime}$ gaining (if $m^{\prime}=m-1$) or losing (if $m^{\prime}=m+1$) a member. This allows to model aggregation and disaggregation mechanisms. Among all the possible processes $e^{\prime}\to e$ (corresponding each to a hyperedge of $\mathcal{E}_{t-1}$ with size $\in\{m-1,m,m+1\}$), the actual one is chosen with probability proportional to the activities $a_{t}(i)$ of $e$’s nodes, to their order propensities $\varphi_{i}(m)$, to the long-term memory $\tilde{\omega}_{0}$ within $e$, and to the duration of the hyperedge $e^{\prime}$. See the Methods for more details. Note that here we consider that only one node can leave/join an existing group at each time step, however more complex splitting and merging mechanisms could potentially be considered [11].
@@ III.3 EATH model parameters
The system parameters include: the average number of interactions $\langle E_{t}\rangle$ and their size distribution $\Psi(m)$, defining the connectivity and the type of group activity; the modulation $\Lambda_{t}$ and the time scales $\mathcal{T},\Delta_{h},\Delta_{l}$, which describe temporal patterns due to specific schedules; the memory matrix heterogeneity $\langle\omega_{0}^{2}\rangle/\langle\omega_{0}\rangle^{2}$ and the probability $p$ that determine the balance between the generation of new hyperedges and the reinforcement of already activated groups. The population properties are: the number of nodes $N$, which defines the size of the population; the distribution of activities $a_{h}(i)$, $a_{T}(i)$, and of the order propensity $\varphi_{i}(m)$, which describe the agents’ heterogeneity. In the next Section we describe how to extract these parameters from a dataset in order to generate a surrogate temporal hypergraph of that dataset.
@@ IV.1 From data to model parameters
First, several parameters are simple quantities that can be straightforwardly extracted from the datasets, such as the population size $N$, the duration $\mathcal{T}$ and the average number of interactions $\langle E_{t}\rangle$.
@@
  • the probability of generating a new hyperedge $p$ is determined by counting at each time-step $t$ the number of hyperedges $O_{t}$ that already existed at time t1t-1 or that have lost/gained one node with respect to time t1t-1:
    $p=\sum_{t=1}^{\mathcal{T}/\delta t}(E_{t}-O_{t})/\sum_{t=1}^{\mathcal{T}/\delta t}E_{t}.$ (8)
@@
In the following subsections, we show the results of the generation of surrogate temporal hypergraphs when using the SFHH dataset as original data [49, 50, 51, 50]: this dataset describes face-to-face interactions occurring during a conference with 403 participants, collected using wearable proximity sensors by the SocioPatterns collaboration over a time span of two days, with a temporal resolution of 20 seconds (see Methods for details). In the SM, we show analogous results for all the other datasets. Moreover, in addition to hypergraphs generated with the EATH model described above, we also consider a baseline version in which no memory effect is taken into account in the generation process (EATHw version): this will allow us to test the importance of the memory mechanisms in reproducing empirical properties.